MAT 152 OL1 Spring 2019 (1) Homework: 3.3 Measures from Grouped Data Score: 0.2 of 1 pt Emely Palacios 1/31/19 7:35 3 of 7 (6 complete) Hw Score: 67.14%, 4.7 3.3.5 EQuestion Help The following data represent the high-temperature distribution for a summer month in a city for some of the last 130 years. Treat the data as a population. Temperature (a) Approximate the mean and standard deviation tor temperature 50-59 2 60-69 314 70-79 1487 80-89 1477 90-99 490 100-109 12 Days H80.3 Round to one decimal place as needed.) ơ 8.3 (Round to one decimal place as needed.) (b) Use the frequency histogram of the data to verify that the distribution is bell shaped 170 No, the frequency histogram of the data is not bel shaped Yes, the frequency histogram of the data is bell shaped 50 110 (c) According to the empirical rule, 95% of days in the month will be between what two temperatures? and (Round to one decimal place as needed. Use ascending order.) Enter your answer in the edit fields and then click Check Answer.

(c) According to the empirical​ rule, 95% of days in the month will be between what two​ temperatures?
Answer #1

Solution:

According to Empirical rule 95% data falls between mu -2*sigma to mu +2*sigma

Hence we need to calculate these first,

μ 80.3

8

Hence ,

mu -2*sigma = 80.3 - 2*16 = 64.3

mu +2*sigma = 80.3 - 2*16 = 96.3

So,

According to the empirical​ rule, 95% of days in the month will be between 64.3 and 96.3

Done